OpenStudy (anonymous):

what are the solutions to the quadratic equation? 2x^2 -16x +32 =0

OpenStudy (paxpolaris):

divide everything by 2

OpenStudy (paxpolaris):

\[x^2-8x+16=0\]

OpenStudy (mayankdevnani):

first of all,understand how to use QUADRATIC FORMULA . http://www.purplemath.com/modules/quadform.htm

OpenStudy (mayankdevnani):

do you know quadratic formula ? @minimoo

OpenStudy (paxpolaris):

\[a^2-2ab+b^2 = (a-b)^2\\ a=x\ ... b=4\] \[(x-4)^2=0 \]

OpenStudy (anonymous):

the -b +- sq rt .....

OpenStudy (skullpatrol):

$$\Huge\text{ It's a perfect square trinomial.}$$

OpenStudy (mayankdevnani):

yeah....somewhat like this !!

OpenStudy (mayankdevnani):